# Neuroevolving Electronic Dynamical Networks

### v1

@@ Abstract
Neuroevolution is a powerful method of applying an evolutionary algorithm to refine the performance of artificial neural networks through natural selection; however, the fitness evaluation of these networks can be time-consuming and computationally expensive, particularly for continuous time recurrent neural networks (CTRNNs) that necessitate the simulation of differential equations. To overcome this challenge, field programmable gate arrays (FPGAs) have emerged as an increasingly popular solution, due to their high performance and low power consumption. Further, their ability to undergo dynamic and partial reconfiguration enables the extremely rapid evaluation of the fitness of CTRNNs, effectively addressing the bottleneck associated with conventional methods. By incorporating fitness evaluation directly upon the programmable logic of the FPGA, hyper-parallel evaluation becomes feasible, dramatically reducing the time required for assessment. This inherent parallelism of FPGAs accelerates the entire neuroevolutionary process by several orders of magnitude, facilitating faster convergence to an optimal solution. The work presented in this study demonstrates the potential of utilizing dynamic and partial reconfiguration on capable FPGAs as a powerful platform for neuroevolving dynamic neural networks.
@@ Introduction
The neuroevolution of Continuous Time Recurrent Neural Networks (CTRNNs) (Hopfield,, 1982; Funahashi and Nakamura,, 1993; Beer,, 1995) is a time-intensive process of evaluating the fitness of time-integrating neural networks, which involves simulating differential equations that govern their dynamics. For non-dynamic neural models, such as convolutional neural networks (LeCun et al.,, 1998; Krizhevsky et al.,, 2012), recurrent neural networks (Bengio et al.,, 1994), long short-term memory networks (Hochreiter and Schmidhuber,, 1997), and many other deep learning methods (Hinton et al.,, 2006; Ng et al.,, 2011), differential calculations are unnecessary during the inference phase and avoided where possible during training, whereas they are required for CTRNNs for both training and inference. In contrast to typical deep learning methods, CTRNNs do not have immediate out-of-the-box support from popular libraries like TensorFlow, PyTorch, or Keras without specific customization, nor do any of these libraries provide hardware acceleration such as Graphical Processing Units (GPUs).
GPUs are currently the standard for training and inference of most deep learning methods due to their ability to perform massive parallel computations, which is well-suited for deep learning’s large matrix operations, resulting in faster training and inference times than traditional CPUs. However, Field Programmable Gate Arrays (FPGAs) can offer a better performance-to-power ratio than GPUs for certain tasks, particularly those that require time-integration. This is because FPGAs can be dynamically reconfigured to execute specific computations efficiently, which is especially useful in tasks that involve expressing complex time-dependent behaviors, including the simulation of dynamical systems or control tasks.
FPGAs offer fine-grained parallelism and reconfigurable hardware, allowing for the implementation of custom-designed circuits tailored to specific computational tasks. This adaptability enables FPGAs to achieve high performance and low latency, making them a suitable choice for real-time applications and edge computing. Furthermore, the low power consumption of FPGAs grants them a competitive edge over GPUs and CPUs, especially in scenarios where energy efficiency is of paramount importance. As the industry evolves, FPGAs are poised to become an increasingly attractive alternative to GPUs, capable of reshaping the landscape of dynamic and liquid artificial neural network training and implementation.
As FPGAs gain traction continue to gain traction, there is a growing interest in harnessing their capabilities for neuroevolution for the following reasons (Floreano et al.,, 2008):
@@
  • Complex fitness landscapes: Neuroevolution is particularly suitable for addressing problems with complex and irregular fitness landscapes, where gradient-based methods might struggle. Evolutionary algorithms are inherently robust and can handle noise, discontinuities, and non differentiability in the fitness function, which may be challenging for gradient descent.
@@
In an effort to unlock each of these benefits, this research explores the use of dynamic and partial reconfiguration of FPGAs to expedite the fitness evaluation of CTRNNs. By integrating the fitness evaluation algorithm into the programmable logic of the FPGA, parallel evaluation becomes feasible, substantially reducing the time required for assessment. In this study, the intricacies of leveraging dynamic and partial reconfiguration of FPGAs as a potent platform for neuroevolving dynamic neural networks is demonstrated, illustrating the potential to overcome the limitations posed by conventional CPU-based evaluation techniques.
@@
In recent years, FPGAs have gained popularity for implementing myriad types of artificial neural networks (ANNs) due to their high performance and low power consumption. Traditionally, ANNs are implemented on general-purpose computing platforms, such as CPUs or GPUs, using software frameworks such as TensorFlow, Keras, or PyTorch. However, as ANNs become larger and more complex, the computational demands of training and running them increase, making them impractical for many applications. FPGAs, on the other hand, offer several advantages for implementing ANNs.
Firstly, FPGAs can be programmed to execute ANNs directly in hardware, which can result in significant performance improvements compared to software-based implementations. This is because FPGAs can perform per-neuron computations in parallel, unlike CPUs and GPUs, which have a limited number of processing cores. In addition, FPGAs can be customized to execute specific operations that are commonly used in ANNs, such as time-dependent dot-product, matrix multiplication, or convolution, further enhancing their performance.
Secondly, FPGAs offer low power consumption, which is particularly important for applications that require ANNs to operate in low-power environments or with limited battery life. This is because FPGAs are designed to allow all components to operate asynchronously and with a low voltage supply, resulting in lower power consumption compared to CPUs and GPUs.
Thirdly, FPGAs offer flexibility and reconfigurability, which can be particularly useful during the development and testing of ANNs. This is because FPGAs can be reconfigured to implement different ANN architectures and parameters, which can enable faster experimentation and prototyping. Moreover, FPGAs can be reconfigured in real-time, allowing for the implementation of hardware-accelerated ANNs that can adapt to changing conditions.
There are several challenges to implementing ANNs on FPGAs, however. One of the main challenges is the design of the ANN architecture and optimization of the FPGA implementation for efficient resource utilization. This requires a deep understanding of both ANNs and FPGAs. Moreover, the programming of FPGAs requires specialized skills and tools, which can be a barrier for some researchers and developers.
@@
Continuous time recurrent neural networks (CTRNNs) are a class of neural networks that are particularly well-suited for modeling dynamical systems. CTRNNs are able to capture the temporal evolution of a system’s state, making them ideal for applications such as control, prediction, and pattern recognition in time-series data(Fernando and Sojakka,, 2003).
At their core, CTRNNs consist of a set of interconnected neurons that evolve over time according to a system of ordinary differential equations. The state of each neuron is represented by a continuous-valued variable, and the connections between neurons are described by a set of weights. The dynamics of the network are governed by the interaction between these weights and the activity of the neurons themselves.
One of the key advantages of CTRNNs is their ability to exhibit complex temporal behaviors. Unlike feedforward neural networks, which operate on a fixed input-output mapping, CTRNNs can generate rich and dynamic output patterns that evolve over time. This makes them particularly useful for tasks such as speech recognition, where the temporal evolution of phonemes plays a critical role in understanding spoken language.
Another strength of CTRNNs is their ability to learn and adapt to new environments. By adjusting the weights of the network connections, CTRNNs can optimize their performance for a given task. This makes them ideal for applications such as robot control, where the ability to learn and adapt in real-time is crucial.
To address these challenges, researchers have developed a range of optimization techniques for training CTRNNs. These include traditional gradient-based methods such as backpropagation through time, as well as more advanced techniques such as evolutionary algorithms and reinforcement learning. These methods allow for the efficient optimization of CTRNNs for a variety of tasks, from prediction and control to classification and generation of complex patterns.
In recent years, the use of CTRNNs has become increasingly widespread, with applications ranging from robotic control to financial forecasting. As the complexity and sophistication of these applications continue to grow, the demand for robust and efficient CTRNN optimization techniques is only set to increase. Through continued research and development, CTRNNs hold the promise of unlocking new levels of performance and adaptability in a range of dynamic systems.
Neuroevolving CTRNNs
CTRNNs are a type of ANN that incorporates time delays in the transmission of signals between neurons, allowing for more complex and dynamic behaviors. However, this also leads to a higher degree of complexity in the fitness evaluation process (Stanley and Miikkulainen,, 2002). The evaluation requires the simulation of differential equations, which can be time-consuming and computationally intensive, especially for large-scale networks. As a result, the bottleneck in evaluating the fitness of CTRNNs through neuroevolution is a significant challenge that needs to be addressed.
A Continuous Time Recurrent Neural Network (CTRNN) is typically defined by a set of differential equations that govern the dynamics of its neurons. Here is an example of the equations for a CTRNN with three neurons:
@@
In these equations, $x_{i}$ represents the state of neuron $i$, and $w_{ij}$ represents the weight of the connection from neuron $j$ to neuron $i$. The inputs $I_{i}$ represent any external inputs to the neurons, and $\sigma$ is an activation function that determines the output of the neuron. The differential equations describe how the state of each neuron changes over time, based on its current state, the inputs it receives, and the weights of its connections to other neurons.
@@
Moreover, evaluating the fitness of CTRNNs in parallel is not straightforward due to the complex nature of the evaluation process. Parallelizing the evaluation of CTRNNs requires careful consideration of the inter-dependencies between neurons and the synchronization of the simulation process. These factors make it challenging to exploit the full potential of parallel computing, further exacerbating the computational time required for fitness evaluation.
Therefore, addressing the bottleneck in fitness evaluation for CTRNNs through neuroevolution is a critical challenge that needs to be addressed to realize the full potential of this optimization technique. Various approaches have been proposed to overcome this challenge, including utilizing hardware acceleration and distributed computing. However, these approaches have their limitations and may not be suitable for all scenarios.
Accelerating Neurovolution with Dynamic and Partial Reconfiguration of FPGAs
@@
As previously described, CTRNNs are an effective class of artificial neural networks that are widely used in a variety of applications such as robotics, control systems, and signal processing. However, evaluating the fitness of CTRNNs trained via neuroevolution has been show to be time-consuming and computationally expensive due to the simulation of differential equations, limiting the scalability of neuroevolutionary approaches for optimizing CTRNNs.
@@
Dynamic and partial reconfiguration of FPGAs allows for the modification of a specific portion of the FPGA hardware architecture while the remaining portion remains active. This capability enables the rapid reconfiguration of the FPGA to perform specific tasks, allowing for parallelism and faster computation time. By implementing the fitness evaluation algorithm on the programmable logic of the FPGA, the evaluation can be performed in parallel, significantly reducing the evaluation time.
@@
The proposed solution offers several advantages over traditional methods. First, the dynamic and partial reconfiguration of FPGAs enables the rapid evaluation of CTRNNs, which can improve the scalability of neuroevolutionary approaches. Second, the intrinsic parallelism of FPGAs can be leveraged to accelerate the neuroevolutionary process, allowing for faster convergence to an optimal solution. Third, the low power consumption of FPGAs compared to traditional processors can result in significant energy savings.
@@
To implement the proposed solution, the fitness evaluation algorithm is programmed onto the FPGA. The CTRNN is then emulated on the FPGA hardware, where the inputs and outputs are stored in the FPGA’s memory. The fitness evaluation function is then performed in parallel by the FPGA for each individual in the population. The evaluated fitness value is then sent back to the host computer for the next generation of the neuroevolution process.
@@
The following sections illustrate the experimental results for the effectiveness of utilizing dynamic and partial reconfiguration of FPGAs for CTRNN optimization through neuroevolution. The proposed method has been shown to significantly reduce the evaluation time of CTRNNs, resulting in a faster convergence to an optimal solution. Additionally, the low power consumption of FPGAs can lead to significant energy savings in large-scale optimization applications.
@@
The purpose of this study is to ascertain the effectiveness of the previously proposed FPGA-based neuroevolution approach. It is important to evaluate its performance and efficiency in comparison to traditional methods, especially CPU-based neuroevolution. This comparative analysis considers various factors, including the time taken for fitness evaluations and the overall computational resources utilized. By demonstrating the superiority of FPGA-based neuroevolution in these aspects, compelling evidence for adopting this method over conventional alternatives is derived.
Another critical aspect of the study is assessing the impact of FPGA-based neuroevolution on the convergence rate of the evolutionary process. Faster convergence to optimal or near-optimal solutions is highly desirable in real-world applications, as it enables more efficient problem-solving and potentially reduces the cost of experimentation. By quantifying the improvement in convergence speed achieved through FPGA-based neuroevolution, future researchers can benefits from this approach.
@@
Lastly, investigating the energy efficiency of the FPGA-based neuroevolution method is of paramount importance, particularly in the context of power consumption and computational resources. As energy efficiency becomes an increasingly critical concern in modern computing, it is essential to demonstrate the advantages of FPGA-based neuroevolution over traditional alternatives in this regard. By quantifying the improvements in energy efficiency achieved through this approach, researchers can further strengthen the case for adopting FPGAs as a powerful platform for neuroevolving CTRNNs and other neural network architectures.
@@
Software as Hardware
In this study, a Xilinx ZCU102 evaluation board (Xilinx,, 2020) was used to embed the fitness evaluation of a CTRNN onto the programmable logic FPGA and subsequently subjected to a differential analysis of the performance of the identical neuroevolutionary task performed on an ARM processor. This required several steps, including the design of the evaluation architecture, communication, and synchronization with the FPGA, as well as optimization of the implementation for performance and resource usage.
@@
To implement the CTRNN model, each PE includes components for neuron state update, activation function computation, and weight matrix multiplication. For simplicity, neuron state updates are computed using the Euler method for numerical integration, as well as a look-up table (LUT) containing 256 entries of fixed-point 16-bit elements describe the standard sigmoid function used to perform these computations. Further, weight matrix multiplication is implemented using pipelined DSP slices. Overall, 628 neurons were able to fit on the ZCU102 without requiring further design optimization.
Communication between the host system and the FPGA is necessary for efficient fitness evaluation. The Xilinx ZCU102 FPGA features a high-speed AXI interface, which can be leveraged to transfer the CTRNN model parameters, including neuron states, connection weights, and time constants, between the FPGA and the onboard ARM processor. The AXI interface can also be used to read the output neuron states from the FPGA after the evaluation is completed. Efficient data transfer is ensured by using Direct Memory Access (DMA) engines, which offload the data movement tasks from the host CPU.
@@
Finally, The Xilinx Internal Configuration Access Port (ICAP) is used to enable dynamic partial reconfiguration (DPR) for the fitness evaluation of CTRNNs on the Xilinx ZCU102 FPGA. DPR allows for the modification of specific regions of the FPGA logic while other parts continue to function without interruption. This capability is essential in scenarios where the fitness evaluation process demands reconfiguration of the processing elements (PEs) or other components within the FPGA fabric to accommodate different CTRNN configurations, activation functions, or optimization techniques. Leveraging the Xilinx ICAP for DPR enables seamless switching between different configurations of the CTRNN evaluation architecture without the need for a full FPGA reconfiguration, which can be time-consuming and disruptive to the overall system operation. By using DPR, multiple CTRNN models with varying complexities can be evaluated in parallel or sequentially, enhancing the overall throughput and efficiency of the system. Furthermore, the ICAP allows for the rapid reconfiguration of the FPGA fabric, enabling adaptive approaches to the fitness evaluation process, such as on-the-fly optimization and, most importantly, updates to CTRNN model parameters at rate of 2.125 Gigaparameters per second .
@@
The results of the comparison between the FPGA and ARM processor-based implementations of the evolutionary algorithm for CTRNNs on the Xilinx ZCU102 platform are presented in this section. The focus is on the total time required to evaluate an increasing number of CTRNNs for an increasing integration duration. The experiments were conducted using a range of time steps per oscillation values (from 100 to 1000, in increments of 100) to determine the performance of both implementations under varying workloads.
Comparison of FPGA and ARM Processor Implementations
The FPGA implementation takes advantage of the inherent parallelism and reconfigurability of the Xilinx ZCU102 platform. The FPGA is able to evaluate a large number of CTRNNs in parallel, which reduces the overall time required to complete the evolutionary process. Furthermore, dynamic and partial reconfiguration allows for the updating of the next generation of CTRNNs without the need for a full reconfiguration of the FPGA, which further reduces the time required for each generation. As a result, the FPGA implementation demonstrates a significant reduction in the total time required to complete 1,000,000 evaluations compared to the ARM processor-based implementation.
@@
The ARM processor implementation, on the other hand, relies on the conventional design of a CPU integrated into the Xilinx ZCU102 platform. And consequently, is unable to match the performance of the FPGA implementation in terms of the total time required to complete the evolutionary process.
Figure 2: Initial comparative performance between FPGA and ARM

### v2

@@ Abstract
Neuroevolution is a powerful method of applying an evolutionary algorithm to refine the performance of artificial neural networks through natural selection; however, the fitness evaluation of these networks can be time-consuming and computationally expensive, particularly for continuous time recurrent neural networks (CTRNNs) that necessitate the simulation of differential equations. To overcome this challenge, methods from evolvable hardware (EHW) offer a solution. The application of evolutionary algorithms to modern system-on-chip (SoC) field programmable gate arrays (FPGAs) that have the ability to undergo dynamic and partial reconfiguration and communicate bilaterally through high speed interfaces (e.g. AXI) enables the extremely rapid evaluation of the fitness of CTRNNs, effectively addressing the update and evaluation bottlenecks associated with conventional methods of evolvable hardware. By incorporating fitness evaluation directly upon the programmable logic of the FPGA, hyper-parallel evaluation becomes feasible, dramatically reducing the time required for assessment. This inherent parallelism of FPGAs accelerates the entire neuroevolutionary process by several orders of magnitude, facilitating faster convergence to an optimal solution. The work presented in this study demonstrates the potential of utilizing FPGAs as a powerful platform for neuroevolving dynamic neural networks.
@@ Introduction
The neuroevolution of Continuous Time Recurrent Neural Networks (CTRNNs) (Hopfield,, 1982; Funahashi and Nakamura,, 1993; Beer,, 1995) is a time-intensive process of evaluating the fitness of time-integrating neural networks, which involves simulating differential equations that govern their dynamics. For non-dynamic neural models, such as convolutional neural networks (LeCun et al.,, 1998; Krizhevsky et al.,, 2012), recurrent neural networks (Bengio et al.,, 1994), long short-term memory networks (Hochreiter and Schmidhuber,, 1997), and many other deep learning methods (Hinton et al.,, 2006; Ng et al.,, 2011), differential calculations are unnecessary during the inference phase and avoided where possible during training, whereas they are required for CTRNNs for both training and inference. In contrast to typical deep learning methods, dynamic neural models such as CTRNNs do not have immediate out-of-the-box support from popular libraries like TensorFlow, PyTorch, or Keras without specific customization, nor do any of these libraries provide hardware acceleration such as general purpose graphical processing units (GPGPUs/GPUs) to dynamic models.
GPGPUs are a type of vector processor currently standard for training most types of deep learning models due to their ability to perform massive parallel computations, which are well-suited for deep learning’s large volume matrix operations, resulting in faster training and inference times than traditional CPUs. However, FPGAs can offer better speed and performance-to-power ratio than GPUs for certain tasks, particularly those that require time-integration. This is because FPGAs can be dynamically reconfigured at the architectural level to execute specific computations efficiently. This ability is critically valuable in tasks that involve expressing complex time-dependent behaviors, such as the simulation of dynamical systems or multivariate control tasks, granting them a competitive edge over GPUs and CPUs, especially in scenarios where energy efficiency is of paramount importance. As the industry evolves, FPGAs are poised to become an increasingly attractive alternative to GPUs, capable of reshaping the landscape of dynamic and liquid artificial neural network training and implementation.
As FPGAs continue to gain traction, there is a growing interest in harnessing their capabilities for neuroevolution for the following reasons (Floreano et al.,, 2008):
@@
  • Complex fitness landscapes: Neuroevolution is particularly suitable for addressing problems with complex and irregular fitness landscapes, where gradient-based methods might struggle. Evolutionary algorithms are inherently robust and can handle noise, discontinuities, and non differentiability in the fitness function, which may be challenging for gradient descent (Amari,, 1993).
@@
By integrating the fitness evaluation algorithm into the programmable logic of the FPGA, parallel evaluation becomes feasible, substantially reducing the time required for assessment. In an effort to unlock each of these benefits, this research explores the use of FPGAs as a potent platform for neuroevolving dynamic neural networks, illustrating the potential to overcome the limitations posed by conventional CPU-based evaluation techniques.
@@
In recent years, FPGAs have gained popularity for implementing myriad types of artificial neural networks (ANNs) due to their high performance and low power consumption. Traditionally, ANNs are implemented on general-purpose computing platforms, such as CPUs or GPUs, using software frameworks such as TensorFlow, Keras, or PyTorch. However, as ANNs become larger and more complex, the computational demands of training and running them increase, making them impractical for many applications. FPGAs, on the other hand, offer several advantages for implementing ANNs:
  1. 1.
    FPGAs can be programmed to execute ANNs directly in hardware, which can result in significant performance improvements compared to software-based implementations. This is because FPGAs can perform per-neuron computations in parallel, unlike CPUs and GPUs, which have a limited number of processing cores. In addition, FPGAs can be customized to execute specific operations that are commonly used in ANNs, such as time-dependent dot-product, matrix multiplication, or convolution, further enhancing their performance.
  2. 2.
    They offer low power consumption, which is particularly important for applications that require ANNs to operate in low-power environments or with limited battery life. This is because FPGAs are designed to allow all components to operate asynchronously and with a low voltage supply, resulting in lower power consumption compared to CPUs and GPUs.
  3. 3.
    FPGAs provide flexibility and reconfigurability, which can be particularly useful during the development and testing of ANNs. This is because they can be reconfigured to implement different ANN architectures and parameters, which can enable faster experimentation and prototyping. Moreover, FPGAs can be reconfigured in real-time, allowing for the implementation of hardware-accelerated ANNs that can adapt to changing conditions.
In light of these advantages, however, one of the main challenges is the design of the ANN architecture and optimization of the FPGA implementation for efficient resource utilization. This requires a deep understanding of both ANNs and FPGAs. Moreover, the programming of FPGAs requires specialized skills and tools, which can be a barrier for some researchers and developers.
@@
Continuous time recurrent neural networks (CTRNNs) are a class of neural networks that are particularly well-suited for modeling dynamical systems (Funahashi and Nakamura,, 1993). CTRNNs are able to capture the temporal evolution of a system’s state, making them ideal for applications such as control, prediction, pattern recognition and generation (Fernando and Sojakka,, 2003).
At their core, CTRNNs consist of a set of interconnected neurons that evolve through time according to a system of ordinary differential equations. The state of each neuron is represented by a continuous-valued variable, and the connections between neurons are described by a set of weights. The overall dynamics of the network are governed by the interaction between these continuous parameters and the activity of the neurons themselves. The model is traditionally written in the following form:
@@
In these equations, $x_{i}$ represents the state of neuron $i$, and $w_{ij}$ represents the weight of the connection from neuron $j$ to neuron $i$. The inputs $I_{i}$ represent any external inputs to the neurons, and $\sigma$ is an activation function (in this case, the standard sigmoid) that determines the output of the neuron. The differential equations describe how the state of each neuron changes over time, based on its current state, the inputs it receives, and the weights of its connections to other neurons.
@@
When compared to most other deep learning models, the key advantage of CTRNNs is their ability to exhibit complex temporal behaviors. Unlike many types of deep learning networks, particularly feedfoward architectures which operate on a fixed input-output mapping, CTRNNs can generate rich and dynamic output patterns that evolve over time. This makes them particularly useful for time series tasks such as speech recognition, gesture recognition, and complex pattern generation, where the time dependent nature of the task domain is integral to the output behavior of the network. Another strength of CTRNNs is their ability to adapt to new environments. By dynamically adjusting the weights of the network, CTRNNs can optimize their performance for a given task. This makes them ideal for applications such as sensing, kinematics, and actuation, where the ability to adapt an output behavior to the environment in real-time is imperative.
To address these challenges, researchers have developed a range of optimization techniques for training CTRNNs. These include traditional gradient-based methods such as backpropagation through time (BPTT), as well as other advanced techniques such as evolutionary algorithms and reinforcement learning. These methods allow for the efficient optimization of CTRNNs for a variety of tasks, from prediction and classification to the generation of complex output patterns and control.
Neuroevolution
Neuroevolution is the method of applying an evolutionary algorithm to refine the parameters of artificial neural networks through an algorithmic process resembling natural selection (Ronald and Schoenauer,, 1994). This approach is quite different from traditional training methods, such as backpropagation, which adjusts the weights of a neural network using gradient descent. Instead, neuroevolution leverages the principles of Darwinian natural selection, mutating and recombining the weights, architectures, or learning rules of neural networks (Stanley and Miikkulainen,, 2002). The selected networks survive and reproduce based on their performance in a given task, effectively allowing evolution to explore vast network configurations and parameter spaces.
Although this method has demonstrated effectiveness in various domains, its primary weakness remains in the computational cost usually associated with the fitness evaluation stage (Xue et al., 2016). Evaluating the fitness of each candidate solution, particularly in the context of CTRNNs which involve the accurate simulation of continuous differential equations, often generates lengthy lead times. This requirement can significantly stretch the duration of evolution, especially when searching for optimal network configurations for complex tasks.
@@
Moreover, evaluating the fitness of CTRNNs in parallel is not straightforward due to the complex nature of the evaluation process. Parallelizing the evaluation of CTRNNs requires careful consideration of the inter-dependencies between neurons and the synchronization of the simulation process. These factors make it challenging to exploit the full potential of parallel computing, further exacerbating the computational time required for fitness evaluation. Therefore, addressing the bottleneck in fitness evaluation for CTRNNs through neuroevolution is a critical challenge that needs to be addressed to realize the full potential of this optimization technique.
Accelerating Neurovolution with FPGAs
@@
As previously described, CTRNNs are an effective class of artificial neural networks that are widely used in a variety of applications such as robotics, control systems, and signal processing. However, evaluating the fitness of CTRNNs trained via neuroevolution is time-consuming and computationally expensive due to the simulation of differential equations, limiting the scalability of neuroevolutionary approaches for optimizing CTRNNs.
@@
Many modern FPGAs have high speed communication interfaces, such as the advanced extensible interface (AXI), and dedicated electronics that enable the modification of a specific portion of the FPGA hardware architecture while the remaining portion remains active (e.g. dynamic partial reconfiguration). This capability enables the rapid reconfiguration of the FPGA to perform specific tasks, allowing for parallelism and faster computation time. By implementing the fitness evaluation algorithm on the programmable logic of the FPGA, the evaluation can be performed in parallel, significantly reducing the evaluation time.
@@
The proposed solution offers several advantages over traditional methods. First, the intrinsic parallelism of FPGAs can be leveraged to accelerate the time series fitness evaluation of CTRNNs, allowing for faster convergence to an optimal solution. Second, FPGAs enables the rapid deployment of CTRNNs to dedicated hardware, which improve the scalability of the neuroevolutionary approach. Third, the low power consumption of FPGAs compared to traditional processors can result in significant energy savings.
@@
To implement the proposed solution, the computational representation of the CTRNN is directly programmed onto the FPGA as a circuit. The CTRNN is then emulated on the FPGA hardware, where the neural parameters are stored in the FPGA’s dedicated block memory. The fitness evaluation function is then performed in parallel by the FPGA for each individual in the population. The network outputs are captured by the host computer and used to calculate the fitness value to be used for the next generation of the neuroevolution process.
@@
The proposed method has been shown to significantly reduce the evaluation time of CTRNNs, resulting in a faster convergence to an optimal solution. Additionally, the low power consumption of FPGAs can lead to significant energy savings in large-scale optimization applications.
@@
The purpose of this study is to determine the effectiveness of the previously proposed FPGA-based neuroevolution approach. It is important to evaluate its performance and efficiency in comparison to traditional methods, especially CPU-based neuroevolution. This comparative analysis considers various factors, including the time taken for fitness evaluations and the overall computational resources utilized. By demonstrating the superiority of FPGA-based neuroevolution in these aspects, compelling evidence for adopting this method over conventional alternatives can be derived.
Another critical aspect of the study is assessing the impact of FPGA-based neuroevolution on the convergence rate of the evolutionary process. Faster convergence to optimal or near-optimal solutions is highly desirable in real-world applications, as it enables more efficient problem-solving and potentially reduces the cost of experimentation. By quantifying the improvement in convergence speed achieved through FPGA-based neuroevolution, future researchers can benefit from this approach.
@@
Lastly, investigating the energy efficiency of the FPGA-based neuroevolution method is important, particularly in the context of power consumption and computational resources. As energy efficiency becomes an increasingly critical concern in modern computing, it is essential to demonstrate the advantages of FPGA-based neuroevolution over traditional alternatives in this regard. By quantifying the improvements in energy efficiency achieved through this approach, researchers can further strengthen the case for adopting FPGAs as a powerful platform for neuroevolving CTRNNs and other neural network architectures.
@@
In this study, a Xilinx ZCU102 evaluation board (Xilinx,, 2020) was used to embed a CTRNN onto the programmable logic FPGA and subsequently subjected to a differential analysis of the performance of the identical neuroevolutionary task performed on an ARM Cortex-A53 processor. This required several steps, including the design of the evaluation architecture, communication, and synchronization with the FPGA, as well as optimization of the implementation for performance and resource usage.
@@
To implement the CTRNN model, each PE includes components for neuron state update, activation function computation, and weight matrix multiplication. For simplicity, neuron state updates are computed using forward Euler method for numerical integration, as well as a look-up table (LUT) containing 256 entries of fixed-point 16-bit elements describing the standard sigmoid function used to efficiently execute these computations. Overall, 628 neurons were able to fit on the ZCU102 without requiring further design optimization.
@@
Communication between the host system and the FPGA is necessary for efficient fitness evaluation. The Xilinx ZCU102 FPGA features a high-speed AXI bus, which is leveraged to transfer the CTRNN model parameters, including neuron states, biases, connection weights, and time constants, between the FPGA and the onboard ARM processor. The AXI interface can also be used to read the output neuron states from the FPGA after the evaluation is completed. Efficient data transfer is ensured by using Direct Memory Access (DMA) engines, which offload the data movement tasks from the host CPU.
The experiment compared the performance of a single ARM core to the performance of an FPGA where each would run an identical copy of a neuroevolutionary algorithm designed to evolve a coupled oscillator. The desired output behavior of the coupled oscillator is such that the output of one neuron would drive the input of the other, and vice versa, generating a sinusoidal-like waveform. The only difference in experimental setup is that the fitness evaluation steps of the evolutionary algorithm were run on the programmable logic of the SoC-based FPGA while the ARM core was used to run the rest of the evolutionary algorithm.
@@
The primary focus for the results is on the total time required to evaluate an increasing number of CTRNNs for an increasing integration duration. The experiments were conducted using a range of evaluation period (from 100 to 1000, in increments of 100) to determine the performance of both implementations under varying workloads.
FPGA versus ARM
The FPGA is able to evaluate a large number of CTRNNs by capitalizing on the intrinsic parallelism of its logic fabric, reducing the overall time required to complete the evolutionary process. Furthermore, the high-bandwidth throughput of the AXI connection allows for the updating of the next generation’s entire population of CTRNNs, which further reduces the time required to evaluate each generation. As a result, the FPGA implementation demonstrates a significant reduction in the total time required to complete the 1,000,000 total evaluations compared to the ARM processor-based implementation.
@@
The ARM processor implementation, on the other hand, which relies on the conventional design of a CPU integrated into the Xilinx ZCU102 platform, is unable to match the performance of the FPGA implementation in terms of the total time required to complete the evolutionary process.
@@
Figure 2: Initial comparative performance between FPGA and ARM